A stone is tossed upward at the moment a ball is dropped from a height of 25 meters. The stone’s initial velocity is 25 m/s. At what height will the two meet? In how many seconds.


1
Expert's answer
2020-10-12T05:27:39-0400

"x_1=25-\frac{gt^2}{2}"


"x_2=v_0t-\frac{gt^2}{2}"


"x_1=x_2\to25=v_0t\to t=25/25=1(s)"


"h=25-9.81\cdot1^2/2=20(m)"


Answer. "h=20(m)" and "t=1(s)" .